OpenStudy (anonymous):

The Patel Manufacturing Company allows for a defect rate of 0.6% on it's production line. The defect average for last week was 4 items out of 4,400. What is the difference between the allowed defect rate and the defect average for last week?

OpenStudy (dumbcow):

4/4400 = 1/1100 = .000999 = .0999%